# Accelerating High-Order Finite Element Simulations at Extreme Scale with FP64 Tensor Cores

### 10 Mar 2026 v1

@@ V Related work
Since the release of tensor cores in NVIDIA’s Volta GPUs [undefa], many HPC applications have successfully leveraged them to improve performance [undefac]. With the addition of FP64 tensor cores in NVIDIA’s Ampere architectures, adoption expanded further, enabling applications requiring double-precision computations to benefit as well [undeff]. Early efforts focused on large matrix operations, primarily using vendor-optimized libraries.
Subsequent FP64 tensor core research has diverged into two paths. For large matrices, the Ozaki scheme [undefad, undefae] emulates FP64 calculations with integer tensor cores, yielding higher performance than FP64 tensor core usage. For smaller matrices, recent studies [undefac, undefj] demonstrate gains from directly employing FP64 tensor cores.
Building on this latter approach, our work extends tensor core acceleration to more complex applications on newer hardware, while also quantifying energy efficiency improvements. Related finite element studies align with these trends. For instance, Cui et al. [undefk] explored NVIDIA A100 tensor cores to speed up tensor-product operations in a finite element Poisson solver. Relevant work was also presented in the CEED-MS40 report [undefaf] of the Center for Efficient Exascale Discretizations [undefag], which examined high-order finite element acceleration using NVIDIA tensor cores and AMD matrix cores for FP64 matrix-matrix multiplications in benchmark kernels on A100 and MI250X GPUs, reporting speedups over vector instructions mainly for orders above 10.
Complementary efforts have leveraged lower-precision tensor cores to balance performance and accuracy in finite element contexts: Ruda et al. [undefah] used reduced-precision tensor cores on NVIDIA A100 and H100 GPUs for matrix-based methods; Yamaguchi et al. [undefai] applied Volta tensor cores to a fast low-order finite element solver for crustal deformation; and Ichimura et al. [undefaj, undefak] developed explicit structured-mesh wavefield simulations with INT8 tensor cores to enhance GPU performance and reduce numerical dispersion.
@@ VI Conclusions
The performance of finite element simulations, including the 2025 Gordon Bell Prize-winning tsunami paper, relies heavily on small matrix operations. In this paper, we showed in detail how to implement and tune such finite element operations using FP64 tensor cores and kernel fusion. By using tensor cores, we reduced shared memory reads by 4.6 and increased the performance of key kernels by up to 59% and energy efficiency by up to 27% on NVIDIA’s GB200 and GH200 chips. Combined with loop fusion, we obtained speedups of up to 2 and energy savings of up to 83%.

### 09 Apr 2026 v2

@@ V Related work
Since the release of tensor cores in NVIDIA’s Volta GPUs [undefa], many HPC applications have successfully leveraged them to improve performance [undefag]. With the addition of FP64 tensor cores in NVIDIA’s Ampere architectures, adoption expanded further, enabling applications requiring double-precision computations to benefit as well [undeff]. Early efforts focused on large matrix operations, primarily using vendor-optimized libraries. Subsequent FP64 tensor core research has diverged into two paths. For large matrices, the Ozaki scheme [undefah, undefai] emulates FP64 calculations with integer tensor cores, yielding higher performance than FP64 tensor core usage. For smaller matrices, recent studies [undefj, undefag] demonstrate gains from directly employing FP64 tensor cores.
Building on this latter approach, our work extends tensor core acceleration to more complex applications on newer hardware (GH200 and GB200), while also quantifying energy efficiency improvements and integrating the optimizations into the production MFEM library for a full PDE-based application. Related finite element studies align with these trends.
For instance, Cui et al. [undefk] explored NVIDIA A100 tensor cores to speed up tensor-product operations in a finite element Poisson solver, reporting significant gains, but did not include kernel fusion or an energy efficiency analysis. Relevant work was also presented in the CEED-MS40 report [undefaj] of the Center for Efficient Exascale Discretizations [undefak], which examined high-order finite element acceleration using NVIDIA tensor cores and AMD matrix cores for FP64 matrix-matrix multiplications in benchmark kernels on A100 and MI250X GPUs, reporting speedups over vector instructions mainly for orders above 10. Our approach builds on these efforts by targeting production-scale, unstructured-mesh operators in a complex tsunami digital twin application, while leveraging MFEM’s modular backends for improved portability across tensor core shapes and precisions.
Complementary efforts have leveraged lower-precision tensor cores to balance performance and accuracy in finite element contexts: Ruda et al. [undefal] used reduced-precision tensor cores on NVIDIA A100 and H100 GPUs for matrix-based methods; Yamaguchi et al. [undefam] applied Volta tensor cores to a fast low-order finite element solver for crustal deformation; and Ichimura et al. [undefan, undefao] developed explicit structured-mesh wavefield simulations with INT8 tensor cores to enhance GPU performance and reduce numerical dispersion.
@@ VI Conclusions
The performance of finite element simulations, including the 2025 Gordon Bell Prize-winning application for real-time tsunami forecasting, relies heavily on small matrix operations. In this paper, we showed in detail how to further improve the performance of this highly-tuned application using FP64 tensor cores and kernel fusion. By using tensor cores, we reduced shared memory reads by 4.6 and increased the performance of key kernels by up to 59% and energy efficiency by up to 27% on NVIDIA’s GB200 and GH200 chips. Combined with loop fusion, we obtained speedups of up to 2 and energy savings of up to 83%.